# Nicer Than Humans: How do Large Language Models Behave in the Prisoner’s Dilemma?

### v1

@@ Abstract
The behavior of Large Language Models (LLMs) as artificial social agents is largely unexplored, and we still lack extensive evidence of how these agents react to simple social stimuli. Testing the behavior of AI agents in classic Game Theory experiments provides a promising theoretical framework for evaluating the norms and values of these agents in archetypal social situations. In this work, we investigate the cooperative behavior of Llama2 when playing the Iterated Prisoner’s Dilemma against random adversaries displaying various levels of hostility. We introduce a systematic methodology to evaluate an LLM’s comprehension of the game’s rules and its capability to parse historical gameplay logs for decision-making. We conducted simulations of games lasting for 100 rounds, and analyzed the LLM’s decisions in terms of dimensions defined in behavioral economics literature. We find that Llama2 tends not to initiate defection but it adopts a cautious approach towards cooperation, sharply shifting towards a behavior that is both forgiving and non-retaliatory only when the opponent reduces its rate of defection below 30%. In comparison to prior research on human participants, Llama2 exhibits a greater inclination towards cooperative behavior. Our systematic approach to the study of LLMs in game theoretical scenarios is a step towards using these simulations to inform practices of LLM auditing and alignment.
@@ 1 Introduction
Early interdisciplinary research has explored the use of LLMs within the context of classical economic games (see Related Work). While highly valuable, all these studies exhibit at least one of the following limitations. First, they generally lack prompt validation procedures, leading to an implicit assumption that LLMs can understand the complex rules of the game and the history of past actions described in the prompt (Akata et al. 2023; Mao et al. 2023; Mei et al. 2024). Second, the duration of simulated games is often limited to a few rounds (Akata et al. 2023; Brookins and DeBacker 2023; Fan et al. 2023; Guo 2023; Xu et al. 2023), hampering the LLMs’ ability to discern the decision-making patterns of other participants – a phenomenon we show in our own experiments. Third, the initialization of LLMs with predefined ‘personas’ tends to skew their responses towards pre-determined behaviors, such as altruism or selfishness (Brookins and DeBacker 2023; Fan et al. 2023; Guo 2023; Horton 2023; Lorè and Heydari 2023; Phelps and Russell 2023). This approach limits the exploration of the LLMs’ baseline behavior, which is crucial for understanding their inherent decision-making processes. Fourth, the evaluation of simulation outcomes has predominantly focused on the quantitative analysis of decision types (e.g., frequency of cooperation), overlooking the LLMs’ higher-level behavioral patterns that can be inferred from the temporal evolution of these decisions (Xu et al. 2023; Mao et al. 2023). The combined effect of these limitations has led to findings that are sometimes inconclusive (Brookins and DeBacker 2023; Mao et al. 2023) and contradictory (Akata et al. 2023; Fan et al. 2023), calling for more systematic evidence on the behavior of LLMs in iterated games.
@@
In this work, we investigate the adaptability of LLMs in terms of their cooperative behavior when facing a spectrum of hostility in an iterated game scenario. We evaluate Llama2 (Touvron et al. 2023) in the Iterated Prisoner’s Dilemma (Osborne and Rubinstein 1994) against adversaries with different propensities for betrayal. Our contribution is threefold. First, we introduce a meta-prompting technique designed to evaluate an LLM’s comprehension of the game’s rules and its ability to parse historical gameplay logs for decision-making. Second, we conduct extensive simulations over 100 rounds and determine the optimal memory span that enables the LLM to adhere to the strategic framework of the game. Third, we analyze the behavioral patterns exhibited by the LLM, aligning them with the core dimensions and strategies delineated in Robert Axelrod’s influential research on the evolution of cooperation within strategic interactions (Axelrod and Hamilton 1981).
We find that Llama2’s behavior is generally ‘nice’, as it tends not to defect first. It adopts a conservative stance towards cooperation, sharply shifting towards behavior that is both forgiving and less retaliatory when the opponent reduces its rate of defection under 30%. Overall, Llama2 displayed a more marked propensity towards cooperation than what existing literature reports about human players.
@@ 2.3 Behavioral Dimensions
Strategies that are $\mathsf{Nice}$, $\mathsf{Forgiving}$, and $\mathsf{Retaliatory}$ (e.g., $\mathsf{TFT}$) perform best against a wide variety of opponents. Human players tend to be particularly uncooperative when faced with games where the reward $R$ for cooperating is much lower than the temptation $T$ to betray the other player. In the indefinitely iterated version of the Prisoner’s Dilemma with a fixed probability at every round for the game to terminate, human subjects tend to be more cooperative when the probability of ending the game is low. Overall, in games with a low chance of continuation and a big gap between $R$ and $T$, the majority of the strategies adopted are most similar to $\mathsf{Always\ Defect}$ (70% to 90%); in games with a longer potential time horizon and $R$ closer to $T$, humans tend to be more forgiving and $\mathsf{Tit\ For\ Tat}$ explains a larger portion of the subjects’ strategies (Dal Bó and Fréchette 2011).
@@ 3.1 LLM Setup
For our experiments, we use the Llama-2-70B-chat model, an open-source LLM, released under a commercial use license [^1] https://ai.meta.com/llama/license/. We use a version of the model hosted by Hugging Face, which we access through the Inference API [^2] https://huggingface.co/inference-api/serverless. We initialize Llama2 with a temperature of $0.7$, in line with the setup of previous work (Lorè and Heydari 2023; Xu et al. 2023).
We make Llama2 play a series of Iterated Prisoner’s Dilemma games, each consisting of $N=100$ rounds. Due to the stochastic nature of the responses that the LLM generates, we repeat each game $k=100$ times and report the average results along with 95% confidence intervals. To evaluate the LLM’s adaptability to different degrees of environmental hostility, we repeat the experiment by matching the LLM against $\mathsf{URND}p$ opponents (defined in §2.2) with varying probability of cooperation $\alpha\in[0,1]$. The final outcome of each game is a sequence containing pairs of binary values representing representing cooperation (1) or defection (0) by the LLM (player $A$) and the opponent (player $B$) at each round $i$:
@@ 3.2 Prompting
To implement the game, we design a fixed system prompt that outlines the game’s rules, including the payoff structure, and the player’s objective to ‘get the highest possible number of points in the long run’. The variable part of the prompt includes the memory of the game, namely a log of the history of the players’ actions up to the current round, along with instructions for generating the next action. The complete prompt can be found in the Appendix. In formulating the memory component, we explore various window sizes to provide the LLM with information from only the $n$ most recent rounds. We determine the optimal window size through targeted experiments that matched the LLM against $\mathsf{AC}$, $\mathsf{AD}$, and $\mathsf{RND}$ players (see §4.2).
@@ 3.3 Meta-prompting
To partially mitigate this issue, we introduce a novel meta-prompting technique to measure the LLM’s comprehension of a given prompt, to inform the process of prompt refinement. Specifically, we formulate a set of prompt comprehension questions that address three key aspects of the prompt (see Table 1): the game rules, to verify the LLM’s grasp of the game mechanics (e.g., ‘What is the lowest payoff that player A can get in a single round?’); the chronological sequence of actions within the game history (e.g., ‘Which action did player A play in round 5?’); and the cumulative game statistics (e.g., ‘What is player’s $B$ current total payoff?’).
To assess the LLM’s proficiency in responding to meta-prompting questions, we conduct three 100-round games against $\mathsf{RND}$ opponents. We pose the questions at each round and compute the average accuracy of the LLM’s responses. At any given round $i$, a question template is instantiated into a set of questions that cover all possible combinations of the template’s parameters. For example, at round $i$, questions referring to specific rounds (actioni and pointsi) are asked for all past rounds from 1 to $i-1$.
@@ 3.4 Behavioral Profiling
We profile the LLM’s behavior with respect to a game history $G^{\alpha}_{k}$ in two ways. First, we quantify behavioral patterns by computing the behavioral dimensions outlined in §2.3. This computation results in a five-dimensional numerical vector that encapsulates the behavioral characteristics of the LLMs. Second, we use the Strategy Frequency Estimation Method (SFEM) defined in previous work (Romero and Rosokha 2018) to calculate the affinity between a player’s game history and any of the classic strategies used in Prisoner Dilemma tournaments (see §2.2). SFEM is a finite-mixture approach that uses likelihood maximization to estimate the likelihood of a strategy being represented in experimental data (Romero and Rosokha 2018). Given a game history, SFEM outputs a score between 00 and $1$ for each strategy in the set of theoretical strategies considered. Being a mixture approach, the sum of all the SFEM scores over the set of possible strategies does not have to sum up to $1$. The final SFEM scores we report are averaged over all the histories analyzed.
@@ 4.1 LLM’s Prompt Comprehension
Figure 1: Accuracy of Llama2 responses to the prompt comprehension questions defined in Table 1. Questions belong to three groups testing respectively aspects related to the rules of the game, its temporal evolution, and its current state. The accuracy of the final full prompt is compared with an earlier version of the prompt that lacked a summary of the cumulative rewards achieved by the players. 95% confidence intervals calculated over 100 rounds are shown.
Figure 1 presents the accuracy of responses to the prompt comprehension questions associated with the final prompt that we used in our experiments, with results averaged over all trial runs. Overall, Llama2 exhibits a good understanding of the concepts assessed by the questions, with most responses achieving an accuracy ranging from $0.8$ to $1.0$.
The process leading to finalizing the prompts was highly iterative, evaluating the accuracy of questions at each incremental step. Generally, adding explicit information about the game state and rules led to a better level or prompt comprehension. For illustration, Figure 1 includes a comparative analysis of accuracy scores derived from an earlier version of the prompt, which lacked a summary of the cumulative point totals for the players. In absence of explicit total counts, Llama2 was required to sum all points from the game history to determine the total tally, which significantly impacted its performance on the #actions and #points questions. This limitation of Llama2 aligns with previous findings that highlight that LLMs tend to struggle with arithmetic (Xu et al. 2023; Aher, Arriaga, and Kalai 2023; Wei et al. 2023). The explicit inclusion of the sum of scores into the prompt markedly improved performance, achieving near-perfect accuracy.
@@ 4.2 Effect of Memory Window Size
Figure 2: Left: Llama2’s probability of cooperation ($p_{coop}$) against an $\mathsf{Always\ Defect}$ opponent, when using a memory window size of 10 vs. including the full game history in the prompt. Right: steady-state probability calculated on the last 10 rounds of $p_{coop}$ for different memory windows sizes. 95% confidence intervals calculated over 100 games are shown.
In iterated games, the information from earlier rounds is essential for a player to deduce the opponent’s strategy, and adapt accordingly. Early research involving LLMs in iterated games experimented with a limited number of rounds, which precluded any analysis of how the size of the memory window influences the agent’s behavior (Akata et al. 2023; Guo 2023; Xu et al. 2023). Here, we analyze the behavior of Llama2 when it is provided with a history of the $n$ most recent rounds. We assess the impact of varying memory window sizes by matching the LLM against an $\mathsf{Always\ Defect}$ opponent. This test relies on the assumption that once the LLM accumulates enough information to infer the consistently defecting behavior of the opponent, the LLM actions should converge towards defection.
Figure 2 (left) shows the probability of the LLM cooperating at each round in games of $100$ rounds, considering two memory window sizes: $m=100$ and $m=10$. Under both conditions, the LLM shifts towards a stance of consistent defection after approximately $5$ to $10$ rounds. Notably, this duration equals or exceeds the maximum number of rounds considered in previous studies. Without any constraints on the history length, the LLM’s cooperation level quickly starts rising back, converging to full cooperation after the $50^{th}$ round. This pattern may be attributed to a combination of two factors: Llama2’s intrinsic preference for positive constructs (Lorè and Heydari 2023) (favoring cooperation over defection) and its limited effectiveness in extracting actionable insights from long prompts (Xi et al. 2023). With a memory window restricted to the $10$ most recent rounds, the LLM retains a full-defection stance throughout the game. We replicated this experiment across a range of memory window sizes and determined their respective equilibrium states by calculating the average cooperation probability in the final $10$ rounds (from the $90^{th}$ to the $100^{th}$). As illustrated in Figure 2 (right), memory windows sizes around $10$ yield the expected outcome. Therefore, we select a window size of $m=10$ for the remainder of our experiments.
@@ 4.3 Behavioral Patterns
Figure 3: Llama2’s probability of cooperation ($p_{coop}$) against $\mathsf{Unfair\ Random}$ opponents with increasing cooperation probability $\alpha$. 95% confidence intervals calculated over 100 games are shown.
We examine the overall propensity for cooperation exhibited by the LLM across various degrees of environmental hostility. Figure 3 shows the relationship between the probability of cooperative behavior $p^{\alpha}_{coop}$ and the varying cooperation levels $\alpha$ of an $\mathsf{Unfair\ Random}$ opponent. The LLM’s inclination to engage in cooperative behavior does not exhibit a linear relationship with $\alpha$. Instead, it follows a sigmoidal curve, indicative of a rapid transition from a predominantly defecting strategy to a more cooperative attitude. The sigmoid is characterized by a relatively flat left tail, maintaining a stable probability of cooperation near $p_{coop}=0$ for $\alpha$ values up to $0.4$. The curve’s inflection point lies between $0.6$ and $0.7$ — well beyond $0.5$ — suggesting that the LLM adopts a cautious approach when interpreting the actions of its opponent, rather than simply mirroring them.
@@
The probability of cooperation provides a macroscopic perspective on a player’s actions. However, to capture more nuanced strategic patterns that emerge during the match, we employed SFEM analysis (defined in §3.4) to estimate the similarity between the behavioral patterns exhibited by the LLM and those commonly observed in games involving human players. Figure 4 presents the SFEM profile for the LLM against increasing values of $\alpha$. When the probability of the adversary’s cooperation surpasses the $0.6$ threshold, there is a discernible shift in the LLM’s strategy from a $\mathsf{Grim\ Trigger}$ strategy to an $\mathsf{Always\ Cooperate}$ approach. The SFEM analysis further indicates that these strategies are the most representative of the LLM’s strategic behavior, with no other strategies being significantly indicative.
At a finer level of analysis, we characterize the behavior of the LLM along the dimensions outlined in Section 2.3 (Figure 5). When the parameter $\alpha$ is set to low values, the LLM exhibits highly uncooperative traits: it frequently retaliates following instances of betrayal, seldom reverts to cooperative behavior after defecting, and is often the initiator of unprovoked defections. As $\alpha$ increases to 0.5 and beyond, we observe a linear transition in these behavioral dimensions, trending towards a more cooperative pattern. The LLM is consistently $\mathsf{Nice}$ across conditions, almost never defecting first. In the scenario where the opponent never defects ($\alpha=1$), the LLM never defects throughout the whole game in $50\%$ of the runs. Interestingly, there is no scenario in which the LLM is simultaneously $\mathsf{Nice}$, $\mathsf{Forgiving}$, and $\mathsf{Retaliatory}$ — the three conditions that characterize the most successful strategies, such as $\mathsf{Tit\ For\ Tat}$.
@@
Figure 4: SFEM scores quantifying the similarity between Llama2’s sequences of actions and known strategies adopted in the Iterated Prisoner’s Dilemma game (defined in §2.2). Llama2’s behavioral sequences come from games against $\mathsf{Unfair\ Random}$ opponents with increasing cooperation probability $\alpha$.
@@
Figure 5: Presence of behavioral traits in Llama2’s actions when playing against $\mathsf{Unfair\ Random}$ opponents with increasing cooperation probability $\alpha$. The values of the same traits calculated for a $\mathsf{Random}$ agent playing against $\mathsf{Unfair\ Random}$ opponents are reported. Some traits are not defined for extreme values of $\alpha$. 95% confidence intervals calculated over 100 games are shown.

### v2

@@ Abstract
The behavior of Large Language Models (LLMs) as artificial social agents is largely unexplored, and we still lack extensive evidence of how these agents react to simple social stimuli. Testing the behavior of AI agents in classic Game Theory experiments provides a promising theoretical framework for evaluating the norms and values of these agents in archetypal social situations. In this work, we investigate the cooperative behavior of three LLMs (Llama2, Llama3, and GPT3.5) when playing the Iterated Prisoner’s Dilemma against random adversaries displaying various levels of hostility. We introduce a systematic methodology to evaluate an LLM’s comprehension of the game rules and its capability to parse historical gameplay logs for decision-making. We conducted simulations of games lasting for 100 rounds and analyzed the LLMs’ decisions in terms of dimensions defined in the behavioral economics literature. We find that all models tend not to initiate defection but act cautiously, favoring cooperation over defection only when the opponent’s defection rate is low. Overall, LLMs behave at least as cooperatively as the typical human player, although our results indicate some substantial differences among models. In particular, Llama2 and GPT3.5 are more cooperative than humans, and especially forgiving and non-retaliatory for opponent defection rates below 30%. More similar to humans, Llama3 exhibits consistently uncooperative and exploitative behavior unless the opponent always cooperates. Our systematic approach to the study of LLMs in game theoretical scenarios is a step towards using these simulations to inform practices of LLM auditing and alignment.
@@ 1 Introduction
Early interdisciplinary research has explored the use of LLMs within the context of classical economic games (see Related Work). While highly valuable, all these studies exhibit at least one of the following limitations. First, they generally lack prompt validation procedures, leading to an implicit assumption that LLMs can understand the complex rules of the game and the history of past actions described in the prompt (Akata et al. 2023; Mao et al. 2023; Mei et al. 2024). Second, the duration of simulated games is often limited to a few rounds (Akata et al. 2023; Brookins and DeBacker 2023; Fan et al. 2023; Guo 2023; Xu et al. 2023), hampering the LLMs’ ability to discern the decision-making patterns of other participants – a phenomenon we show in our own experiments. Third, the initialization of LLMs with predefined ‘personas’ tends to skew their responses towards pre-determined behaviors, such as altruism or selfishness (Brookins and DeBacker 2023; Fan et al. 2023; Guo 2023; Horton 2023; Lorè and Heydari 2023; Phelps and Russell 2023). This approach limits the exploration of the LLMs’ baseline behavior, which is crucial for understanding their inherent decision-making processes. Last, the evaluation of simulation outcomes has predominantly focused on the quantitative analysis of decision types (e.g., frequency of cooperation), overlooking the LLMs’ higher-level behavioral patterns that can be inferred from the temporal evolution of these decisions (Xu et al. 2023; Mao et al. 2023). The combined effect of these limitations has led to findings that are sometimes inconclusive (Brookins and DeBacker 2023; Mao et al. 2023) and contradictory (Akata et al. 2023; Fan et al. 2023), calling for more systematic evidence on the behavior of LLMs in iterated games.
@@
In this work, we investigate the adaptability of LLMs in terms of their cooperative behavior when facing a spectrum of hostility in an iterated game scenario. We evaluate Llama2 (Touvron et al. 2023), Llama3, and GPT3.5 (Brown et al. 2020) in the Iterated Prisoner’s Dilemma (Osborne and Rubinstein 1994) against adversaries with different propensities for betrayal. Our contribution is threefold. First, we introduce a meta-prompting technique designed to evaluate an LLM’s comprehension of the game’s rules and its ability to parse historical gameplay logs for decision-making. Second, we conduct extensive simulations over 100 rounds and determine the optimal memory span that enables the LLMs to adhere to the strategic framework of the game. Third, we analyze the behavioral patterns exhibited by the LLMs, aligning them with the core dimensions and strategies delineated in Robert Axelrod’s influential research on the evolution of cooperation within strategic interactions (Axelrod and Hamilton 1981).
We observe that, overall, the three models tend to be more cooperative than humans, but they display some variability in their strategies. This variability persists even when the models are exposed to the same environment, game, and task framing. Both Llama2 and GPT3.5 displayed a more marked propensity towards cooperation than what existing literature reports about human players, indicating a favorable alignment with positive values. In contrast, Llama3 adopts a more strategic and exploitative approach that is more similar to that of humans. This approach may be advantageous in competitive scenarios where raw performance is critical, but it can be a disadvantage when it comes to aligning with positive values.
@@ 2.3 Behavioral Dimensions
Strategies that are $\mathsf{Nice}$, $\mathsf{Forgiving}$, and $\mathsf{Retaliatory}$ (e.g., $\mathsf{TFT}$) perform best against a wide variety of opponents. Human players tend to be particularly uncooperative when faced with games where the reward $R$ for cooperating is much lower than the temptation $T$ to betray the other player. In the indefinitely iterated version of the Prisoner’s Dilemma with a fixed probability at every round for the game to terminate, human subjects tend to be more cooperative when the probability of ending the game is low. Usually, in games with a low chance of continuation and a big gap between $R$ and $T$, the majority of the strategies adopted are most similar to $\mathsf{Always\ Defect}$ (70% to 90%); in games with a longer potential time horizon and $R$ closer to $T$, humans tend to be more forgiving and $\mathsf{Tit\ For\ Tat}$ explains a larger portion of the subjects’ strategies (Dal Bó and Fréchette 2011).
@@ 3.1 LLM Setup
In our experiments, we use Llama-2-70b-chat-hf and Llama-3-70B-Instruct as open-source language models developed by Meta and released under commercial use licenses [^1] https://ai.meta.com/llama/license/, [^2] https://llama.meta.com/llama3/license/. We access these models through the Hugging Face platform, using their Inference API [^3] https://huggingface.co/inference-api/serverless. As a closed-source language model, we use GPT-3.5-Turbo, developed and hosted by OpenAI, accessed via their proprietary API [^4] https://openai.com/index/openai-api/. GPT3.5 has been used in many early experiments on LLM agents in game-theoretical scenarios (Mei et al. 2024; Lorè and Heydari 2023; Xu et al. 2023).
The models are initialized with a temperature value of $0.7$, that is equal to the default value for GPT models and is consistent with previous studies using models from the Llama family (Lorè and Heydari 2023; Xu et al. 2023). An analysis of robustness to different temperature settings ($0.1$ and $1.0$) is provided in the Appendix (Figure A9).
We make the LLMs play a series of Iterated Prisoner’s Dilemma games, each consisting of $N=100$ rounds. Due to the stochastic nature of the responses that LLMs generate, we repeat each game $k=100$ times and report the average results along with 95% confidence intervals. To evaluate the models’ adaptability to different degrees of environmental hostility, we repeat the experiment by matching them against $\mathsf{URND}p$ opponents (defined in §2.2) with varying probability of cooperation $\alpha\in[0,1]$. The final outcome of each game is a sequence containing pairs of binary values representing the actions of the LLM (player $A$) and the opponent (player $B$) at each round $i$:
@@ 3.2 Prompting
To implement the game, we design a fixed system prompt that outlines the game’s rules, including the payoff structure, and the player’s objective to ‘get the highest possible number of points in the long run’. The variable part of the prompt includes the memory of the game, namely a log of the history of the players’ actions up to the current round, along with instructions for generating the next action. The complete prompt can be found in the Appendix (Figures A1, A2, and A3). In iterated games, the information from earlier rounds is essential for a player to deduce the opponent’s strategy, and adapt accordingly. Early research involving LLMs in iterated games experimented with a limited number of rounds, precluding any analysis of how the size of the memory window influences the agent’s behavior (Akata et al. 2023; Guo 2023; Xu et al. 2023). In formulating the memory component, we explore various window sizes to provide the model with information from only the $n$ most recent rounds. We evaluate the effect of different memory window sizes by testing the LLM against an $\mathsf{Always\ Defect}$ opponent and identifying the optimal window size (see §4.2). This assessment is based on the premise that, once the LLM has gathered sufficient information to recognize the opponent’s consistently defecting behavior, its actions should align with defection, which is the only logical strategy.
@@ 3.3 Meta-prompting
To partially mitigate this issue, we introduce a novel meta-prompting technique to measure the LLMs’ comprehension of a given prompt, to inform the process of prompt refinement. Specifically, we formulate a set of prompt comprehension questions that address three key aspects of the prompt (see Table 1): the game rules, to verify the LLM’s grasp of the game mechanics (e.g., ‘What is the lowest payoff that player A can get in a single round?’); the chronological sequence of actions within the game history (e.g., ‘Which action did player A play in round 5?’); and the cumulative game statistics (e.g., ‘What is player’s $B$ current total payoff?’). To assess the LLMs’ proficiency in responding to meta-prompting questions, we conduct a series of $3$ games of $100$ rounds each against $\mathsf{RND}$ opponents. We pose the questions at each round and compute the average accuracy of the LLMs’ responses. At any given round $i$, a question template is instantiated into a set of questions that cover all possible combinations of the template’s parameters. For example, at round $i$, questions referring to specific rounds (actioni and pointsi) are asked for all past rounds from 1 to $i-1$.
@@ 3.4 Behavioral Profiling
We profile the LLMs’ behavior with respect to a game history $G^{\alpha}_{k}$ in two ways. First, we quantify behavioral patterns by computing the behavioral dimensions outlined in §2.3. This computation results in a five-dimensional numerical vector that encapsulates the behavioral characteristics of the LLMs. Second, we use the Strategy Frequency Estimation Method (SFEM) defined in previous work (Romero and Rosokha 2018) to calculate the affinity between a player’s game history and any of the classic strategies used in Prisoner Dilemma tournaments (see §2.2). SFEM is a finite-mixture approach that uses likelihood maximization to estimate the likelihood of a strategy being represented in experimental data (Romero and Rosokha 2018). Given a game history, SFEM outputs a score between 00 and $1$ for each strategy in the set of theoretical strategies considered. Being a mixture approach, the sum of all the SFEM scores over the set of possible strategies does not have to sum up to $1$. The final SFEM scores we report are averaged over all the histories analyzed.
@@ 4.1 LLM’s Prompt Comprehension
Figure 1: Accuracy of the models’ responses to the prompt comprehension questions defined in Table 1. The questions are categorized into three groups, each assessing different aspects: the rules of the game, its temporal evolution, and its current state. We show 95% confidence intervals, computed from 100 games.
Figure 1 presents the accuracy of responses to the prompt comprehension questions associated with the final prompt that we used in our experiments, with results averaged over all trial runs. Overall, the models exhibit a good understanding of the concepts assessed by the questions, with most responses achieving an accuracy ranging from $0.8$ to $1.0$.
We iteratively tested multiple prompt versions against the response accuracy of Llama2. No further iterations of other models were needed, as the best prompt for Llama2 yielded very high response accuracy in both Llama3 and GPT3.5. Generally, adding explicit information about the game state and rules led to a better level of prompt comprehension. For illustration, Figure A7 in the Appendix includes a comparative analysis of accuracy scores derived from an earlier version of the prompt, which lacked a summary of the cumulative point totals for the players. In the absence of explicit total counts, Llama2 was required to sum all points from the game history to determine the total tally, which significantly impacted its performance on the #actions and #points questions. This limitation aligns with previous findings that highlight that LLMs tend to struggle with arithmetic (Xu et al. 2023; Aher, Arriaga, and Kalai 2023; Wei et al. 2023). The explicit inclusion of the sum of scores into the prompt markedly improved performance, achieving near-perfect accuracy.
@@ 4.2 Effect of Memory Window Size
Figure 2: Left: Llama2’s probability of cooperation ($p_{coop}$) against an $\mathsf{Always\ Defect}$ opponent, when using a memory window size of 10 vs. including the full game history in the prompt. Right: steady-state probability calculated on the last 10 rounds of $p_{coop}$ for different memory windows sizes. We show 95% confidence intervals, computed from 100 games.
Figure 2 (left) shows the probability of the LLM cooperating at each round in games of $100$ rounds, considering two memory window sizes: $m=100$ and $m=10$. Under both conditions, the LLM shifts towards a stance of consistent defection after approximately $5$ to $10$ rounds. Notably, this duration equals or exceeds the maximum number of rounds considered in previous studies. Without any constraints on the history length, the LLM’s cooperation level quickly starts rising back, converging to full cooperation after the $50^{th}$ round. This pattern may be attributed to a combination of two factors: Llama2’s intrinsic preference for positive constructs (Lorè and Heydari 2023) (favoring cooperation over defection) and its limited effectiveness in extracting actionable insights from long prompts (Xi et al. 2023). With a memory window restricted to the $10$ most recent rounds, the LLM retains a full-defection stance throughout the game. We replicated this experiment across a range of memory window sizes and determined their respective equilibrium states by calculating the average cooperation probability in the final $10$ rounds (from the $90^{th}$ to the $100^{th}$). As illustrated in Figure 2 (right), memory windows sizes around $10$ yield the expected outcome. Therefore, we select a window size of $m=10$ for the remainder of our experiments. Differently, Llama3 and GPT3.5 show no variation in their cooperation levels regardless of the memory window size (see Figure A8 in Appendix), allowing us to use the same prompt designed for Llama2. Notably, Llama3 does not suffer from the same bias towards cooperation as Llama2, and keeps its defective behavior also when provided with the full game history. GPT3.5 exhibits the less optimal behavior, converging towards defection much more slowly.
@@ 4.3 Behavioral Patterns
Figure 3: Models’ probability of cooperation ($p_{coop}$) against $\mathsf{Unfair\ Random}$ opponents with increasing cooperation probability $\alpha$. We show 95% confidence intervals, computed from 100 games.
Probability of cooperation
We examine the overall propensity for cooperation exhibited by the LLM across various degrees of environmental hostility. Figure 3 shows the relationship between the probability of cooperative behavior $p^{\alpha}_{coop}$ of each model and the varying cooperation levels $\alpha$ of an $\mathsf{Unfair\ Random}$ opponent.
Among the three models, Llama3 demonstrates the most strategic approach. It maintains a very low level of cooperation even when the opponent is nearly always cooperating ($p_{coop}<0.3$ for every $\alpha<1$), but it increases its cooperation to nearly 100% when the opponent is $\mathsf{AC}$ ($\alpha=1$). In contrast, Llama2’s behavior follows a sigmoidal curve across the entire range of $\alpha$ from 00 to $1$. This indicates a rapid transition from a predominantly defecting strategy to a more cooperative attitude. The sigmoid curve is characterized by a relatively flat left tail, maintaining a stable probability of cooperation near $p_{coop}=0$ for $\alpha$ values up to $0.4$. The curve’s inflection point is between $0.6$ and $0.7$—well beyond $0.5$—suggesting a cautious approach in interpreting the opponent’s actions. Compared to Llama3, the older Meta model is less guarded and more prone to increase its cooperation as soon as the opponent’s cooperation significantly rises over the 50/50 chance. GPT3.5 displays a less strategic approach, maintaining a low but still significant cooperation level ($0.40.2<p_{coop}<$ 0.4) for low $\alpha$ ($<0.5$) and not reaching full cooperation even when the opponent is $\mathsf{AC}$.
In general, the models exhibit non-linear behaviors but with different characteristics. A common trend is the increasing cooperation as $\alpha$ grows, demonstrating a minimum level of strategic behavior for all LLMs.
These findings, especially concerning Llama3, support previous research that highlights the cautious response patterns of LLMs in repeated game scenarios (Akata et al. 2023; Phelps and Russell 2023).
SFEM profile
@@
The probability of cooperation provides a macroscopic perspective on a player’s actions. However, to capture more nuanced strategic patterns that emerge during the match, we employed SFEM analysis (defined in §3.4) to estimate the similarity between the behavioral patterns exhibited by the LLMs and those commonly observed in games involving human players.
Figure 4 illustrates which strategies best explain the behaviors of each model as the values of $\alpha$ increase. When the adversary’s probability of cooperation exceeds the $0.6$ threshold, there is a noticeable shift in both Llama2’s and GPT3.5’s strategy from $\mathsf{Grim\ Trigger}$ to $\mathsf{Always\ Cooperate}$. In contrast, Llama3 consistently aligns with the $\mathsf{Grim\ Trigger}$ strategy across all values of $\alpha$, displaying an exploitative approach even when the opponent tends to cooperate for the majority of the time. SFEM analysis further indicates that these strategies are the most representative of the LLMs’ strategic behavior, with no other strategies being significantly indicative.
Previous studies have shown that humans tend to be particularly uncooperative when the reward $R$ is much lower than the temptation $T$ and when playing for shorter time horizons, typically opting for startegies close to $\mathsf{Always\ Defect}$ (Dal Bó and Fréchette 2011; Romero and Rosokha 2018). Conversely, humans become more cooperative when the reward is much higher and the time horizon is longer, opting for a mixture of $\mathsf{Tit\ For\ Tat}$ and $\mathsf{Grim\ Trigger}$ strategies.
Comparing this prior knowledge to our results indicates that GPT3.5 and Llama2 are consistently more cooperative than humans. In situations that do not favor cooperation—where the opponent defects more frequently than cooperates—they adopt the $\mathsf{GRIM}$ strategy, unlike humans who tend to use the $\mathsf{AD}$ strategy. Similarly, when the context supports cooperation (i.e., the opponent cooperates more than half of the time), both GPT3.5 and Llama2 employ the $\mathsf{AC}$ strategy, which is notably more cooperative than the human preference for $\mathsf{TFT}$ or $\mathsf{GRIM}$. On the other hand, Llama3 is more cooperative than humans in hostile environments $(p_{\text{coop}}<0.5$), also using the $\mathsf{GRIM}$ strategy. However, in environments that encourage cooperation, Llama3 behaves more like humans by continuing to use the $\mathsf{GRIM}$ strategy, whereas humans typically mix between the more forgiving $\mathsf{TFT}$ strategy and $\mathsf{GRIM}$.
@@
Figure 4: SFEM scores quantifying the similarity between the models’ sequences of actions and known strategies adopted in the Iterated Prisoner’s Dilemma game (defined in §2.2). The models’ behavioral sequences come from games against $\mathsf{Unfair\ Random}$ opponents with increasing cooperation probability $\alpha$. Some SFEM scores are not shown because not well-defined for extreme values of $\alpha$.
Behavioral profile
At a finer level of analysis, we characterize the behavior of the LLMs along the dimensions outlined in Section 2.3 (Figure 5).
When the parameter $\alpha$ is set to low values, the models exhibit highly uncooperative traits: they frequently retaliate following instances of betrayal, seldom revert to cooperative behavior after defecting, and are often the initiators of unprovoked defections. GPT3.5 exhibits these traits in a more moderate manner, whereas Llama2 and Llama3 display more extreme profiles in this regard. More noticeable differences among the models’ response emerge beyond $\alpha=0.5$. While GPT3.5 and Llama2 generate choices that are more $\mathsf{Forgiving}$ and less $\mathsf{Troublemaking}$, Llama3 remains $\mathsf{Troublemaking}$ and rarely $\mathsf{Forgiving}$ for every $\alpha<1.0$. On the other hand, the three models are constantly $\mathsf{Nice}$ across conditions, rarely defecting first. GPT3.5 shows the lowest scores but remains above 0.5 even against an opponent that never defects ($\alpha=1$). In this scenario, all the LLMs never defect throughout the whole game in $50\%$ of the runs. Interestingly, there is no scenario in which any LLM is simultaneously $\mathsf{Nice}$, $\mathsf{Forgiving}$, and $\mathsf{Retaliatory}$ — the three conditions that characterize the most successful strategies, such as $\mathsf{Tit\ For\ Tat}$.
@@
Figure 5: Presence of behavioral traits in the models’ actions when playing against $\mathsf{Unfair\ Random}$ opponents with increasing cooperation probability $\alpha$. The values of the same traits calculated for a $\mathsf{Random}$ agent playing against $\mathsf{Unfair\ Random}$ opponents are reported. Some traits are not defined for extreme values of $\alpha$. We show 95% confidence intervals, computed from 100 games.